# Oracle-based Uniform Sampling from Convex Bodies

### v1

@@ 3 Projection oracle-based proximal sampling
Assume conditions (A1), (A2) and the step size $\eta=1/d^{2}$. Consider Algorithm 2 with the RGO implementation via a projection oracle and rejection sampling (Algorithm 3). Let $\epsilon>0$. Denote $C_{\mathrm{PI}}$ and $C_{\mathrm{LSI}}$ respectively the Poincaré constant and the log Sobolev constant of the uniform distribution $\pi^{X}$ on $K$, whose asymptotics are provided in Lemma 2.1.
  1. a)
    To reach $\epsilon$-accuracy in Rényi divergence $\mathcal{R}_{q}$, the ASF (i.e., Algorithm 2) takes at most $O\left(d^{2}{C_{\mathrm{LSI}}q\log\left(\frac{2\log M}{\epsilon}\right)}\right)$ iterations. In each iteration, Algorithm 3 as the RGO implementation makes one query to the projection oracle on $K$. Moreover, the average number of rejections for the rejection sampling in each iteration is no more than $M(\sqrt{2\pi e}+1)$.
  2. b)
    To reach $\epsilon$-accuracy in $\chi^{2}$-divergence, the ASF takes at most $O\left(d^{2}{C_{\mathrm{PI}}\log\left(2\frac{M^{2}+1}{\epsilon}\right)}\right)$ iterations. In each iteration, the number of projection oracle queries and average number of rejection for the rejection sampling are the same as those in Part a.
4 Separation oracle-based proximal sampling
In Section 3, under the assumption that a projection oracle is available, one gets an exact solution to $\mathrm{argmin}\,_{x\in\mathbb{R}^{d}}\Theta^{\eta,K}_{y}(x)$ as $\operatorname{proj}_{K}(y)$, and the Gaussian proposal for the rejection sampling can thus be centered at $\operatorname{proj}_{K}(y)$. However, what happens if a projection oracle is not available? In that scenarios, we propose to use a state-of-the-art Cutting Plane method by [27], which uses a separation oracle on $K$ to find an approximate solution $\hat{x}$ of $\mathrm{argmin}\,_{x\in\mathbb{R}^{d}}\Theta^{\eta,K}_{y}(x)$. From there, one can once again implement rejection sampling to complete the RGO step.
After finding a suitable $\hat{x}$, we still need to construct a proposal for the rejection sampling. The proposal we choose is
@@ 4 Separation oracle-based proximal sampling
Compared to the Gaussian proposal $\mathcal{N}\left(\operatorname{proj}_{K}(y),\eta I_{d}\right)$ in Section 3, one can see the above proposal follows a similar spirit in the sense that it is centered at a high-concentration point $\hat{x}$. It should be noted that $\nu(x)$ is no longer a Gaussian distribution. Generating a sample for this non-Gaussian proposal turns out to be simple as it can be transformed into a one-dimensional sampling problem, as Lemma B.3 in the Appendix B will show.
Below are our implementation of the RGO under the assumption that we have access to a separation oracle on $K$. As before, ${\cal U}[0,1]$ denotes the uniform distribution on $[0,1]$.
Algorithm 4 Implementation of the RGO with separation oracle
1. Generate a $(1/d)$-solution $\hat{x}$ of $\mathrm{argmin}\,_{x\in\mathbb{R}^{d}}\Theta^{\eta,K}_{y}(x)$ using the Cutting Plane method by [27], which employs a separation oracle on $K$.
2. Via Algorithm 5 in Appendix B.4, generate $X\sim\nu(x)$. Also, generate $U\sim{\cal U}[0,1]$.
3. If
$U\leq\frac{\exp\left(-\Theta_{y}^{\eta,K}(X)\right)}{\exp\left(-\mathcal{P}_{2}(X)\right)},$ (16)
then accept $X$; otherwise, reject $X$ and go to step 2. The function $\mathcal{P}_{2}$ is defined in (17).
In particular, Lemma B.2 in Appendix B.1 provides the number of separation oracle calls for the Cutting Plane method to generate $\hat{x}$ in Step 1, while Lemma A.3 in Appendix A states that the sample $X$ generated by Algorithm 4 follows the distribution $\pi^{X|Y}$, thus ensuring that the rejection sampling is unbiased.
The following result is analogous to Lemma 3.2. The acceptance test at (16) is well-defined only if the right hand side of (16) is no more than 1, which means we need to show $\Theta^{\eta,K}_{y}(x)\geq\mathcal{P}_{2}(x),\forall x$. The idea is to show $\mathcal{P}_{1}\geq\mathcal{P}_{2}$ (where $\mathcal{P}_{1}$ is defined in (13)) and combine it with Lemma 3.2 which says $\Theta_{y}^{\eta,K}\geq\mathcal{P}_{1}$. The proof of the next result is deferred to Appendix B.2.
Lemma 4.1
Define
Recall $\Theta^{\eta,K}_{y}$ and $\mathcal{P}_{1}$ defined in (9) and (13), respectively. Then, we have for every $x\in\mathbb{R}^{d}$,
$\Theta^{\eta,K}_{y}(x)\geq\mathcal{P}_{1}(x)\geq\mathcal{P}_{2}(x).$
In particular, the fact that $\Theta^{\eta,K}_{y}\geq\mathcal{P}_{2}$ ensures the acceptance test at (16) is well-defined.
Next, we are able to deduce the average number of rejections in Algorithm 4 for each iteration of the ASF (Algorithm 2). The proof of the upcoming result is deferred to Appendix B.3.
Proposition 4.2
Assume conditions (A1), (A2), and the step size $\eta=1/d^{2}$. Then the average number of rejections in Algorithm 4 is no more than $\sqrt{2\pi}M\exp\left(\frac{13}{4}+\frac{20}{d}\right)+M\exp\left(\frac{9}{4}+\frac{12}{d}\right)$.
The following theorem presents the main result of Section 4. We omit the proof as it immediately follows after putting together Corollary 2.3 for the iteration complexity of the outer loops of the ASF (Algorithm 2), Proposition 4.2 for the average number of rejections of an inner loop, and Lemma B.2 for the number of separation oracle queries of an inner loop.
Theorem 4.3
Assume conditions (A1), (A2), and the step size $\eta=1/d^{2}$. Consider Algorithm 2 with the implementation via a separation oracle and rejection sampling (Algorithm 4). Denote $C_{\mathrm{PI}}$ and $C_{\mathrm{LSI}}$ respectively the Poincaré constant and the log Sobolev constant of the uniform distribution $\pi^{X}$ on $K$ whose asymptotics are provided in Lemma 2.1.
  1. a)
    To reach $\epsilon$-accuracy in Rényi divergence $\mathcal{R}_{q}$, the ASF, i.e., Algorithm 2, takes at most $O\left(d^{2}{C_{\mathrm{LSI}}q\log\left(\frac{2\log M}{\epsilon}\right)}\right)$ iterations. In each iteration, Algorithm 4 as the RGO implementation makes ${\cal O}\left(d\log\frac{d\gamma}{\alpha}\right)$ queries to the separation oracle on $K$, where
    $\gamma=\frac{R}{\mbox{minwidth}(K)},\qquad\mbox{minwidth}(K)=\min_{\left\lVert a\right\rVert=1}\left(\max_{y\in K}a^{T}y-\min_{y\in K}a^{T}y\right),$ (18)
    and $\alpha\in(0,1)$ satisfies the concentration inequality
    $\Pr\left(\alpha\leq\frac{2}{d^{3}7R^{2}}\right)\leq 4\exp\left(-\frac{d^{2}R^{2}}{8}\right).$
    Moreover, the average number of rejections for the rejection sampling in each iteration is no more than $\sqrt{2\pi}M\exp\left(\frac{13}{4}+\frac{20}{d}\right)+M\exp\left(\frac{9}{4}+\frac{12}{d}\right)$.
  2. b)
    To reach $\epsilon$-accuracy in $\chi^{2}$-divergence, the ASF takes at most $O\left(d^{2}{C_{\mathrm{PI}}\log\left(2\frac{M^{2}+1}{\epsilon}\right)}\right)$ iterations. Per an iteration of the ASF, the number of separation oracle queries and the average number of rejections for the rejection sampling are the same as those in Part a.
@@ 5 Concluding remarks
In this paper, we propose algorithms for uniform sampling from a convex body $K$ based on the ASF/proximal sampler. We explore the use of either the projection oracle on $K$ or the separation oracle on $K$ for the RGO implementation (Algorithm 3 and Algorithm 4, respectively). Our RGO implementations are exact and therefore our algorithms do not have any failure probability. In both cases, the algorithms perform ${\cal O}(d^{2})$ RGO steps. With a projection oracle, each RGO queries one projection and has at most ${\cal O}(1)$ expected rejections. With a separation oracle, each RGO queries ${\cal O}(d\log d)$ separations and has at most ${\cal O}(1)$ expected rejections.
We finally discuss some possible extensions of the paper. First, a natural question to ask beyond uniform sampling on $K$ is general log-concave sampling on $K$. Both uniform sampling on $K$ and log-concave sampling on $\mathbb{R}^{d}$ have benefited from using ASF as a generic framework in recent years; as a consequence, it is interesting to investigate algorithms based on ASF for sampling $\exp(-f(x))$ on $K$. Second, for the purpose of uniform sampling on $K$, the RGO implementations in this paper (i.e., Algorithm 3 and Algorithm 4) and those in [39, 37, 34] all require a small step size $\eta=1/d^{2}$ so that RGO implementations within ASF remain efficient. In contrast, for sampling from $\exp(-f(x))$ on $\mathbb{R}^{d}$, under the assumption that $f$ satisfies an $(L_{\alpha},\alpha)$-semi-smooth condition for some $\alpha\in[0,1]$, the step size condition can be relaxed to $\eta=\tilde{\cal O}(d^{-\alpha/(\alpha+1)})$ in [18]. In particular, this improves the dimension dependence from ${\cal O}(d)$ to ${\cal O}(\sqrt{d})$ when $f$ is smooth (i.e., $\alpha=1$). However, techniques in [18] cannot be directly applied to uniform sampling on $K$, since the log-density (i.e., the indicator function $I_{K}(x)$) is discontinuous and hence lacks a smoothness notion. Therefore, reducing the dimension dependence for uniform sampling on $K$ still remains a challenging yet meaningful question.
@@ Acknowledgement
  • Appendix A contains some technical lemmas and short proofs.
  • Appendix B contains results that are relevant to Algorithm 4, for example the Cutting Plane method by [27].
  • Appendix C provides a brief summary of some algorithms for uniform sampling from convex bodies.
Appendix A Technical lemmas and proofs
@@
Lemma A.2
Denote $\partial K$ the boundary set of $K$ and assume condition (A1) holds. Then, we have
@@
Next, we provide the proof of Corollary 2.3 based on the contraction result in Theorem 2.2.
@@
Proof of Corollary 2.3: The $M$-warm start assumption implies $\mathcal{R}_{q}\left(\mu^{X}_{0}||\pi^{X}\right)\leq\frac{q}{q-1}\log M$. Then via the first part of (4) in Theorem 2.2, we can solve for

### v2

@@ 4 Separation oracle-based proximal sampling
Assume conditions (A1), (A2), and the stepsize $\eta=1/d^{2}$. Then regarding Algorithm 4,
  • a)
    the average number of proposals in Algorithm 4 is no more than $\sqrt{2\pi}M\exp\left(\frac{13}{4}+\frac{20}{d}\right)+M\exp\left(\frac{9}{4}+\frac{12}{d}\right)$.
  • b)
    there are at most ${\cal O}\left(d\log\frac{d\gamma}{\alpha}\right)$ queries to the separation oracle on $K$, where $\gamma=\frac{R}{\mbox{minwidth}(K)}$ and $\alpha\in(0,1)$ satisfies the concentration inequality $\Pr\left(\alpha\leq\frac{2}{7d^{3}R^{2}}\right)\leq 4\exp\left(-\frac{d^{2}R^{2}}{8}\right)$.
5 Numerical experiments on dense $Z$-polytopes
We compare our PS+Alg3 (Algorithm 2+Algorithm 3) with In-and-Out (Algorithm 1 by [41]) on on random dense Z-polytopes. Per the explanation given in Appendix D, a Z-polytope is specified by a generator matrix $V=[v_{1},\dots,v_{m}]\in\mathbb{R}^{d\times m}$ through
$Z\;=\;\{x=Vt:\ \|t\|_{\infty}\leq 1\}\;=\;\left\{\sum_{j=1}^{m}t_{j}v_{j}:\ t_{j}\in[-1,1]\right\}.$
Setup.
We draw $V\in\mathbb{R}^{d\times m}$ with i.i.d. standard Gaussian entries and normalize each column to unit $\ell_{2}$ norm for numerical stability. We set $d=40$ and $m=80$, and run $n=10^{5}$ outer iterations per seed over $6$ independent seeds. Following Theorem 23 in [41], we take $M=1$ and failure probability $p=0.1$, which yields the conservative In-and-Out stepsize $\eta_{\mathrm{inout}}=\left(2d^{2}\log\!\Bigl(\frac{9nM}{p}\Bigr)\right)^{-1}\approx 1.95\times 10^{-5}.$
We will consider two stepsizes:
$\eta_{A}:=\eta_{\mathrm{inout}}\,\text{(Exp. A, conservative)}\quad\text{and}\quad\eta_{B}:=10^{-2}\,\text{(Exp. B, aggressive)}.$
Note that the dominant scaling in $\eta_{A}$ is $1/d^{2}$; while the dependence on $n,M,p$ enters only through the logarithmic factor. In Experiment A, we run both samplers with the conservative stepsize $\eta_{A}$ that adheres to the recommendation of our Theorem 3.5 as well as [41, Theorem 23] . In Experiment B, we rerun both samplers at the larger practical $\eta_{B}$ to stress-test performance outside the conservative theory. A baseline reference marginal is computed once using a long hit-and-run chain. All experiments are done in Matlab with the Optimization Toolbox. The starting point of PS+Alg3 and In-and-Out is the central symmetry point of the Z-polytope.
For Z-polytopes, $x\in\mathbb{R}^{d}$ is represented as $x=Vt$ with generator coefficients $t\in[-1,1]^{m}$. As explained in Appendix D, both the membership check and projection for Z-polytopes are implemented via interior-point methods and have comparable arithmetic costs. Specifically, membership of $x$ is checked by LP feasibility (find $t\in[-1,1]^{m}$ with $Vt=x$, using linprog), and projection $\Pi_{Z}(y)$ is computed by the box-constrained QP $\min_{t\in[-1,1]^{m}}\|Vt-y\|_{2}^{2}$ (using quadprog) and returning $Vt^{\star}$. We report the elapsed time/wall-clock time, including all oracle-solver overhead.
(a) Exp. A ($\eta_{A}$): TV vs. iterations
(b) Exp. A ($\eta_{A}$): TV vs. elapsed time
(c) Exp. B ($\eta_{B}$): TV vs. iterations
(d) Exp. B ($\eta_{B}$): TV vs. elapsed time
Figure 1: Comparison of In-and-Out and PS+Alg3.
Results.
@@
Experiment A (conservative stepsize $\eta_{A}$). Figure 1(a) plots the TV error of the $x_{1}$ marginal versus outer iterations under the matched conservative stepsize $\eta_{A}$. From Figure 1(a), after about $5\times 10^{4}$ iterations, PS+Alg3 reduces TV faster and reaches a substantially lower TV level than In-and-Out. Meanwhile, Figure 1(b) shows TV versus elapsed time, where elapsed time captures both algorithmic complexity and oracle-solver overhead. The figure shows PS+Alg3 is slower in elapsed time, but it reaches a substantially lower TV level and thus has a better accuracy-time tradeoff.
Experiment B (aggressive stepsize $\eta_{B}$). Figure 1(c) reports TV versus outer iterations of both samplers at the larger stepsize $\eta_{B}=10^{-2}$. In this instance, In-and-Out remains stable at $\eta_{B}$ and its TV curve continues to decrease, reaching a final TV level comparable to PS+Alg3. PS+Alg3 still shows a modest advantage in early iterations and attains a comparable final TV. Figure 1(d) presents TV versus elapsed time in the same regime: PS+Alg3 reaches a given TV substantially sooner in wall-clock time, whereas In-and-Out requires longer elapsed time to achieve similar accuracy.
@@ 6 Concluding remarks
In this paper, we propose algorithms for uniform sampling from a convex body $K$ based on the proximal sampler. We explore the use of either the projection oracle on $K$ or the separation oracle on $K$, standard tools in convex optimization, for the RGO implementation (Algorithm 3 and Algorithm 4, respectively). Our RGO implementations are exact and therefore our algorithms do not have any failure probability. In both cases, the algorithms perform ${\cal O}(d^{2})$ RGO steps. With a projection oracle, each RGO queries one projection and has at most ${\cal O}(1)$ expected proposals. With a separation oracle, each RGO queries ${\cal O}(d\log d)$ separations and has at most ${\cal O}(1)$ expected proposals. In addition, we perform numerical experiments to compare In-and-Out (Algorithm 1 by [41]) with PS+Alg3 (Algorithm 2+Algorithm 3) on random dense $Z$-polytopes. In the experiments of Section 5, PS+Alg3 outperforms In-and-Out, both at the conservative stepsize suggested by the theory and at a more aggressive stepsize, even though the two methods invoke different oracles (projection versus membership).
We finally discuss some possible extensions of the paper. First, a natural question to ask beyond uniform sampling on $K$ is general log-concave sampling on $K$. Both uniform sampling on $K$ and log-concave sampling on $\mathbb{R}^{d}$ have benefited from using the proximal sampler as a generic framework in recent years; as a consequence, it is interesting to investigate algorithms based on the proximal sampler for sampling $\exp(-f(x))$ on $K$. Second, for the purpose of uniform sampling on $K$, the RGO implementations in this paper (i.e., Algorithm 3 and Algorithm 4) and those in [41, 43, 45] all require a small stepsize $\eta=1/d^{2}$ so that RGO implementations within the proximal sampler remain efficient. In contrast, for sampling from $\exp(-f(x))$ on $\mathbb{R}^{d}$, under the assumption that $f$ satisfies an $(L_{\alpha},\alpha)$-semi-smooth condition for some $\alpha\in[0,1]$, the stepsize condition can be relaxed to $\eta=\tilde{\cal O}(d^{-\alpha/(\alpha+1)})$ in [24]. In particular, this improves the dimension dependence from ${\cal O}(d)$ to ${\cal O}(\sqrt{d})$ when $f$ is smooth (i.e., $\alpha=1$). However, techniques by [24] cannot be directly applied to uniform sampling on $K$, since the negative log-density (i.e., the indicator function $I_{K}(x)$) is discontinuous and hence lacks a smoothness notion. Therefore, reducing the dimension dependence for uniform sampling on $K$ still remains a challenging yet meaningful question.
@@ Acknowledgement
  • Appendix A provides a summary of some algorithms for uniform sampling from convex bodies.
  • Appendix B collects some technical lemmas and short proofs deferred from Sections 2 and 3.
  • Appendix C contains results that are relevant to Algorithm 4, for example, the Cutting Plane method by [33].
  • Appendix D presents examples of common convex bodies and their oracle implementations.
Appendix A Algorithms for uniform sampling on convex bodies
  • The Hit-and-Run walk is first introduced in [67] and rigorously investigated by Lovász and Simonovits in [59]. Also assuming a membership oracle on $K$ and in the special case of uniform sampling, the Hit-and-Run walk is: choose a uniform direction over the unit sphere and find a line segment in that direction that intersects $K$ at two endpoints but still belong to $K$; then go to a uniform random point on that line segment. [61] shows the its iteration complexity in total variance is of the order $\mathcal{O}\left(d^{2}C_{\mathrm{LSI}}\frac{M^{2}}{\epsilon^{2}}\right)$.
  • The In-and-Out algorithm by [41] is the proximal sampler by [48], where the RGO implementation is via a membership oracle: sample $x$ from a Gaussian distribution until $x\in K$ up to a certain number of maximum attempts $N$, at which point the algorithm halts and declares failure. Their iteration complexities in Rényi divergence and $\chi^{2}$-divergence are the same as those in our Theorem 2.2 in Section 2, except that they derive their iteration complexities via the PI constant of the uniform distribution while we use both PI and LSI constants.
Appendix B Technical results and proofs
@@
Lemma B.2
Denote $\partial K$ the boundary set of a non-empty, closed, and convex set $K\in\mathbb{R}^{d}$ such that $B(0,1)\subseteq K$. Then, we have
@@
In Algorithm 3 and Algorithm 4, we use a simple rejection-sampling scheme: propose from a tractable density proportional to $\exp(-\mathcal{P})$, and accept with the standard likelihood-ratio rule in order to target a density proportional to $\exp(-\Theta)$. In the upcomoing lemma, we record the resulting target law, the acceptance probability, and the average number of proposals.
Lemma B.3
Let $\Theta,\mathcal{P}:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ be measurable and assume
$0<N_{\Theta}:=\int_{\mathbb{R}^{m}}e^{-\Theta(z)}\mathrm{d}z<\infty,\qquad 0<N_{\mathcal{P}}:=\int_{\mathbb{R}^{m}}e^{-\mathcal{P}(z)}\mathrm{d}z<\infty,$
and that $\Theta(z)\geq\mathcal{P}(z)$ for all $z\in\mathbb{R}^{m}$. Consider the rejection sampler that repeats the following trial until acceptance:
  1. 1.
    draw a proposal $Z\in\mathbb{R}^{m}$ with density $k(z)=e^{-\mathcal{P}(z)}/N_{\mathcal{P}}$;
  2. 2.
    draw $U\sim{\cal U}[0,1]$ independent of $Z$ and accept if and only if
    $U\leq\exp\bigl(-\Theta(Z)+\mathcal{P}(Z)\bigr).$ (15)
Then the output has density
$\pi(z)=\frac{e^{-\Theta(z)}}{N_{\Theta}}.$
Moreover, let $E$ denote the acceptance event (15) for a single trial and $F$ denote the number of trials until acceptance. Then, the acceptance probability and the average number of proposals are
$p:=\Pr(E)=\frac{N_{\Theta}}{N_{\mathcal{P}}},\qquad\mathbb{E}[F]=\frac{1}{p}=\frac{N_{\mathcal{P}}}{N_{\Theta}}.$ (16)
Proof: For one trial, let $E$ denote the event (15) happens. Since $\Theta\geq\mathcal{P}$, we have $0\leq\Pr(E\mid Z=z)=\exp(-\Theta(z)+\mathcal{P}(z))\leq 1$, so the acceptance rule is well-defined.
We first compute the acceptance probability:
$p=\Pr(E)=\int_{\mathbb{R}^{m}}\Pr(E\mid Z=z)k(z)\mathrm{d}z$ $=\int_{\mathbb{R}^{m}}\exp\bigl(-\Theta(z)+\mathcal{P}(z)\bigr)\cdot\frac{e^{-\mathcal{P}(z)}}{N_{\mathcal{P}}}\mathrm{d}z$
$=\frac{1}{N_{\mathcal{P}}}\int_{\mathbb{R}^{m}}e^{-\Theta(z)}\mathrm{d}z=\frac{N_{\Theta}}{N_{\mathcal{P}}}.$
Next, we compute the conditional density of the accepted proposal:
$k(z\mid E)$ $=\frac{\Pr(E\mid Z=z)k(z)}{\Pr(E)}=\frac{\exp\bigl(-\Theta(z)+\mathcal{P}(z)\bigr)\cdot e^{-\mathcal{P}(z)}/N_{\mathcal{P}}}{N_{\Theta}/N_{\mathcal{P}}}=\frac{e^{-\Theta(z)}}{N_{\Theta}}.$
Thus the accepted proposal has density $\pi(z)\propto e^{-\Theta(z)}$, as claimed.
Finally, each trial is i.i.d. and succeeds with probability $p$, hence $F\sim\mathrm{Geom}(p)$ and $\mathbb{E}[F]=1/p$.
Next we record a straightforward implication of the warm start assumption (A2).
Lemma B.4
Let $\mu^{X}_{k}$ denote the law of $x_{k}$ and $\mu^{k}$ denote the law of $y_{k}$ in Algorithm 2. Under the warm start condition (A2), the warmness is preserved for both the $x$ and $y$-updates at every iteration $k\geq 0$:
$\frac{\mathrm{d}\mu^{k}}{\mathrm{d}\pi^{Y}}\leq M\quad\text{and}\quad\frac{\mathrm{d}\mu^{X}_{k}}{\mathrm{d}\pi^{X}}\leq M.$
Proof: We proceed by induction. The base case $k=0$ holds by assumption (A2). Assume as the induction hypothesis that $\frac{\mathrm{d}\mu^{X}_{k}}{\mathrm{d}\pi^{X}}\leq M$.
Step 1: Recall $\mu^{k}$ is the distribution of $y_{k}$ obtained in Step 1 of Algorithm 2. For any measurable set $U\subseteq\mathbb{R}^{d}$, let $U-y:=\{x\in\mathbb{R}^{d}:x+y\in U\}$. Denote $\gamma(\cdot)$ the density of $\mathcal{N}(0,\eta I_{d})$. Since $\mu^{k}=\mu^{X}_{k}*\gamma$ and $\pi^{Y}=\pi^{X}*\gamma$ where * denotes the convolution of the probability measures, we have
$\mu^{k}(U)=\int_{\mathbb{R}^{d}}\mu^{X}_{k}(U-y)\gamma(y)\mathrm{d}y$ $=\int_{\mathbb{R}^{d}}\left(\int_{U-y}\frac{\mathrm{d}\mu^{X}_{k}}{\mathrm{d}\pi^{X}}(t)\pi^{X}(\mathrm{d}t)\right)\gamma(y)\mathrm{d}y$
$\leq M\int_{\mathbb{R}^{d}}\pi^{X}(U-y)\gamma(y)\mathrm{d}y=M\pi^{Y}(U),$
where we have applied the induction hypothesis to get the last line. This implies $\frac{\mathrm{d}\mu^{k}}{\mathrm{d}\pi^{Y}}\leq M$.
Step 2: Let $\mu^{X}_{k+1}$ denote the distribution of $x_{k+1}$ obtained in Step 2 of Algorithm 2. Recall that $x_{k+1}$ is sampled from the conditional distribution $\pi^{X|Y}(\cdot|y_{k})$. For any measurable set $A\subseteq K$, we have
$\mu^{X}_{k+1}(A)$ $=\int_{\mathbb{R}^{d}}\pi^{X|Y}(A|y)\mu^{k}(\mathrm{d}y)$
$=\int_{\mathbb{R}^{d}}\pi^{X|Y}(A|y)\frac{\mathrm{d}\mu^{k}}{\mathrm{d}\pi^{Y}}(y)\pi^{Y}(\mathrm{d}y)$
$\leq\int_{\mathbb{R}^{d}}\pi^{X|Y}(A|y)M\pi^{Y}(\mathrm{d}y)=M\int_{\mathbb{R}^{d}}\pi^{X,Y}(A,\mathrm{d}y)=M\pi^{X}(A),$
where we apply $\frac{\mathrm{d}\mu^{k}}{\mathrm{d}\pi^{Y}}\leq M$ from Step 1 to get the second to last line. Thus, $\frac{\mathrm{d}\mu^{X}_{k+1}}{\mathrm{d}\pi^{X}}\leq M$. The induction proof is complete.
Recall that Lemma 3.3 shows the following relation
$\int_{K^{c}}\exp\left(-\frac{\left(\|\operatorname{proj}_{K}(y)-y\|-\tau\right)^{2}}{2\eta}\right)\mathrm{d}y=\int_{0}^{\infty}\exp\left(-\frac{\left(\delta-\tau\right)^{2}}{2\eta}\right)\mathsf{vol}_{d-1}(\partial K_{\delta})\mathrm{d}\delta,$
and provides an upper bound on it. The following result presents an alternative upper bound.
Lemma B.5
Assume the setup in Lemma 3.3 and write $v(\delta)=\mathrm{vol}(K_{\delta})$ and $a(\delta)=\mathrm{vol}_{d-1}(\partial K_{\delta})$. Then, $v^{\prime}(\delta)=a(\delta)$ for almost every $\delta\geq 0$. Fix $\eta>0$ and $\tau\geq 0$, and set
$w_{\eta,\tau}(\delta)=\exp\!\left(-\frac{(\delta-\tau)^{2}}{2\eta}\right).$
It holds that
$\int_{0}^{\infty}a(\delta)w_{\eta,\tau}(\delta)\mathrm{d}\delta\leq\mathsf{vol}(K)\Bigl(e^{d\tau}-1\Bigr)+\mathsf{vol}(K)\sqrt{2\pi\eta d^{2}}\;\exp\!\left(\tau d+\frac{\eta d^{2}}{2}\right).$
Proof: By the co-area formula, we have $v(\delta)=\mathrm{vol}(K_{\delta})=\int_{0}^{\delta}\mathrm{vol}_{d-1}(\partial K_{s})\mathrm{d}s=\int_{0}^{\delta}a(s)\mathrm{d}s$, so that $v^{\prime}(\delta)=a(\delta)$ for almost every $\delta\geq 0$. Then integrating by parts gives
$\int_{0}^{\infty}a(\delta)w_{\eta,\tau}(\delta)\mathrm{d}\delta=\int_{0}^{\infty}v^{\prime}(\delta)w_{\eta,\tau}(\delta)\mathrm{d}\delta=v(\delta)w_{\eta,\tau}(\delta)\bigg|_{0}^{\infty}-\int_{0}^{\infty}v(\delta)w_{\eta,\tau}^{\prime}(\delta)\mathrm{d}\delta.$ (17)
Noting that $v(\delta)w_{\eta,\tau}(\delta)\to 0$ as $\delta\to\infty$, we have
$\int_{0}^{\infty}a(\delta)w_{\eta,\tau}(\delta)\mathrm{d}\delta=-\mathrm{vol}(K)\exp\!\left(-\frac{\tau^{2}}{2\eta}\right)+\int_{0}^{\infty}v(\delta)\frac{\delta-\tau}{\eta}w_{\eta,\tau}(\delta)\mathrm{d}\delta.$ (18)
It follows from the observation $K\subseteq K_{\delta}\subseteq(1+\delta)K$ that
$\mathsf{vol}(K)\leq v(\delta)\leq(1+\delta)^{d}\mathrm{vol}(K)\leq e^{d\delta}\mathrm{vol}(K).$ (19)
For $\delta\in[0,\tau]$, we have
[…]
where the last identity uses $w^{\prime}_{\eta,\tau}(\delta)=-\frac{\delta-\tau}{\eta}w_{\eta,\tau}(\delta)$. For $\delta\in[\tau,\infty)$, using (19), we obtain
$\int_{\tau}^{\infty}v(\delta)\frac{\delta-\tau}{\eta}w_{\eta,\tau}(\delta)\mathrm{d}\delta\leq\mathsf{vol}(K)\int_{\tau}^{\infty}e^{d\delta}\frac{\delta-\tau}{\eta}w_{\eta,\tau}(\delta)\mathrm{d}\delta.$
Plugging the above two inequalities into (18) yields
$\int_{0}^{\infty}a(\delta)w_{\eta,\tau}(\delta)\mathrm{d}\delta=-\mathsf{vol}(K)+\mathsf{vol}(K)\int_{\tau}^{\infty}e^{d\delta}\frac{\delta-\tau}{\eta}w_{\eta,\tau}(\delta)\mathrm{d}\delta.$ (20)
Let $F(\delta):=e^{d\delta}w_{\eta,\tau}(\delta)$, then we note that
$\frac{\delta-\tau}{\eta}F(\delta)=dF(\delta)-F^{\prime}(\delta).$
It follows from $F(\delta)\to 0$ as $\delta\to\infty$ and $F(\tau)=e^{d\tau}$ that
$\int_{\tau}^{\infty}e^{d\delta}\frac{\delta-\tau}{\eta}w_{\eta,\tau}(\delta)\mathrm{d}\delta$ $=\int_{\tau}^{\infty}\Bigl(dF(\delta)-F^{\prime}(\delta)\Bigr)\mathrm{d}\delta$
$=d\int_{\tau}^{\infty}e^{d\delta}w_{\eta,\tau}(\delta)\mathrm{d}\delta-F(\delta)\bigg|_{\tau}^{\infty}=d\int_{\tau}^{\infty}e^{d\delta}w_{\eta,\tau}(\delta)\mathrm{d}\delta+e^{d\tau}.$ (21)
Plugging (21) into (20) yields
$\int_{0}^{\infty}a(\delta)w_{\eta,\tau}(\delta)\mathrm{d}\delta\leq\mathsf{vol}(K)\Bigl(e^{d\tau}-1\Bigr)+d\mathsf{vol}(K)\int_{\tau}^{\infty}e^{d\delta}w_{\eta,\tau}(\delta)\mathrm{d}\delta.$ (22)
Finally, we complete the square by writing $d\delta-\frac{(\delta-\tau)^{2}}{2\eta}=-\frac{\left(\delta-(\tau+\eta d)\right)^{2}}{2\eta}+\tau d+\frac{\eta d^{2}}{2}$, so
$\int_{\tau}^{\infty}e^{d\delta}w_{\eta,\tau}(\delta)\mathrm{d}\delta$ $\leq\int_{-\infty}^{\infty}\exp\!\left(d\delta-\frac{(\delta-\tau)^{2}}{2\eta}\right)\mathrm{d}\delta$
$=\exp\!\left(\tau d+\frac{\eta d^{2}}{2}\right)\int_{-\infty}^{\infty}\exp\!\left(-\frac{\left(\delta-(\tau+\eta d)\right)^{2}}{2\eta}\right)\mathrm{d}\delta$
$=\sqrt{2\pi\eta}\;\exp\!\left(\tau d+\frac{\eta d^{2}}{2}\right),$ (23)
where the last identity follows from Lemma B.1(a). Combining this with (22) gives, for every $\tau\geq 0$, the desired bound on $\int_{0}^{\infty}a(\delta)w_{\eta,\tau}(\delta)\mathrm{d}\delta$. This completes the proof.
We provide two missing proofs in Section 2 and Section 3 below.
Proof of Lemma 3.2 It follows from the convexity of $K$ (condition (A1)) that
$\langle x-\operatorname{proj}_{K}(y)\,,\operatorname{proj}_{K}(y)-y\rangle\geq 0,$
and hence that the first stated inequality in Lemma 3.2 holds. In view of (4) and (6), we observe that the RHS of (5) is equivalent to
Hence, the proof is completed.
At this point, we recall an important result by [41] about contractivity in Rényi and $\chi^{2}$ divergences for Algorithm 2. Using a clever smoothing argument, [41] adapts the approach of [11] to uniform sampling over a convex body $K$. Moreover, as noted in [41], the result does not actually require $K$ to be convex.
Theorem B.6
([41, Theorem 23]) Let $\mu^{X}_{k}$ be the law of the $k$-th output of Algorithm 2. Denote $C_{\mathrm{PI}}$ and $C_{\mathrm{LSI}}$ respectively the Poincaré constant and the log Sobolev constant of the uniform distribution $\pi^{X}$ on $K$ whose asymptotics are provided in Lemma 2.1. Then for any $q\geq 1$,
Now, we are ready to prove Theorem 2.2 based on the contraction result in Theorem B.6.
@@
Proof of Theorem 2.2: The $M$-warm start assumption implies $\mathcal{R}_{q}\left(\mu^{X}_{0}||\pi^{X}\right)\leq\frac{q}{q-1}\log M$. Then via the first part of (24) in Theorem B.6, we can solve for